# Gravitational Lensing of Gravitational Waves from Astrophysical Sources: Theory, Detection, and Applications

### 07 May 2026 v1

@@ 1 Introduction
Detection of the gravitational wave (GW) signals from inspiralling and mergers of compact objects [including black holes (BHs) and neutron stars (NSs)] provide a new way to measure their properties and study their origin and nature. The first direct detection of a stellar binary black hole (sBBH) merger (GW150914) by the Laser Interferometer Gravitational wave Observatory (LIGO) confirmed the existence of sBBHs and substantially revised our understanding of sBBH formation (e.g., Abbott et al. 2016a, 2019a, 2020b; The LIGO Scientific Collaboration et al. 2021a, b). The subsequent observations of the binary neutron star (BNS) merger event, GW170817, including its multiwavelength electromagnetic (EM) counterparts not only confirmed the predictions about short $\gamma$-ray burst (sGRB) with afterglow signals and kilonova produced by BNS mergers, but also marked the opening of the new era of multi-messenger astronomy (e.g., Abbott et al. 2017a, b; Coulter et al. 2017; Pian et al. 2017; Dobie et al. 2018; Coughlin et al. 2019; Abbott et al. 2019b). With the multimessenger (GW + EM) information, one may estimate the mass of the tidal debris of BNS by the observation of the kilonova signals and therefore further constrain the fundamental physics related to the equation of state (EOS) of neutron stars (NSs) (e.g., Most et al. 2018; De et al. 2018; Gill et al. 2019). Moreover, with the redshift measurement from the EM signals and the luminosity distance inferred from GW, one can also constrain the Hubble constant to a precision of $\sim 10\%$ with this single object (GW170817) by the “standard siren” method (e.g., Schutz 1986; Abbott et al. 2017c; Soares-Santos et al. 2019). These results underscore the strong prospects for observing compact-binary coalescences (CBCs) via GWs and EMs.
@@
The astrophysical GW sources are quite abundant in the universe and an extremely large number of such sources are expected to be detected in future. Since the first direct detection of GWs by LIGO, more than $300$ stellar CBCs have been detected by LIGO and Virgo, including about three hundreds sBBH mergers, two BNS mergers, and several BHNS mergers (e.g., The LIGO Scientific Collaboration et al. 2025b, c)111see also at the website https://gracedb.ligo.org/. With further upgrade, the Advanced LIGO plus (LIGO A+) (Abbott et al. 2020a) and LIGO Voyager (Adhikari et al. 2020) will detect many more in the near future. The third-generation ground-based GW detectors, i.e., Einstein Telescope (ET) (Hild et al. 2011) and Cosmic Explorer (CE) (Reitze et al. 2019) are even anticipated to detect more than several tens of thousands of CBCs per year. In addition, the pulsar timing arrays (PTA) (e.g., Agazie et al. 2023a; EPTA Collaboration et al. 2023; Reardon et al. 2023; Xu et al. 2023; Miles et al. 2025) are expected to detect the nanohertz GW signals from inspiralling supermassive binary black holes (SMBBHs) with masses $\gtrsim 10^{8}M_{\odot}$ (e.g., Sesana 2013a; Chen et al. 2024a) both collectively as a stochastic background and individually as continuous gravitational wave sources (CGWs). The space-based GW detectors, such as Laser Interferometer Space Antenna (LISA) (Amaro-Seoane et al. 2017), Taiji (Luo et al. 2020) and Tianqin (Mei et al. 2021), are designed to detect the millihertz GW signals from mergers of massive binary black holes (MBBHs) with masses $\sim 10^{4}-10^{7}M_{\odot}$ and extreme/intermediate mass-ratio inspirals (EMRIs/IMRIs) (e.g., Amaro-Seoane et al. 2007; Babak et al. 2017). The numbers of such sources in the millihertz and nanohertz bands that may be detected in future can also be substantial.
@@
A small fraction of various types of GW events may be gravitationally lensed by intervening objects/systems, such as stars, galaxies, and galaxy clusters, and they may be detectable and identifiable by future GW detectors. Since the GWs considered for detection (by ground-based GW detectors, space-based GW detectors, and the PTAs) have much longer wavelengths ($\lambda_{\rm GW}$) comparing with those of EM waves from radio to gamma-ray bands, the wave nature of GWs may have to be considered in some cases, especially when $\lambda_{\rm GW}\gg\lambda_{\rm E}$ with $\lambda_{\rm E}$ representing the Einstein radius of the lens (a characteristic scale). In those cases with $\lambda_{\rm GW}\ll\lambda_{\rm E}$, the gravitational lensing of GWs is in the geometric-optics regime, and the lensing effect is similar to that for EM waves (such as lensed QSOs, lensed galaxies, lensed planets, etc.). However, in some cases the GW wavelength can be larger than the Einstein radius of the lens, i.e., $\lambda_{\rm GW}\gtrsim\lambda_{\rm E}$, thus the wave-optics should apply for the lensing and the diffraction is important, which hardly occur for the cases of EM wave lensing. In addition, not like the cases for EM waves that may be easily contaminated by light from foreground galaxies or distorted by interstellar medium (ISM), the GW signals are hardly contaminated and thus encode clean information about physical properties of the sources and the propagation path. Therefore, the lensing of GW signals offers unique a probe for some key astrophysical and cosmological questions (e.g., Grespan & Biesiada 2023; Smith et al. 2025; More & Phurailatpam 2025; Levan et al. 2025), including the wave nature of GWs, the nature of the dark matter (e.g., Guo & Lu 2022; Fairbairn et al. 2023), the origin of sBBH mergers (e.g., Chen et al. 2022b), and the cosmological parameters at relatively high redshifts (Liao et al. 2017; Hannuksela et al. 2020; Chen et al. 2025a; Shan et al. 2023).
@@ 2.1.1 GW Waveform
The GW waveform is different in different evolution stages of binary mergers, i.e., inspiral, merger, and ring-down. In the merger and ring-down stages, the waveform can only be accurately calculated by using numerical relativity, for their extremely complex spacetime structure, which is computationally expensive. In the inspiral stage, one can analytically obtain the waveform, and such a method also enables an order of magnitude estimation for the waveform in the other two stages (Bondi 1960; Bondi et al. 1962). Combining the analytical and numerical methods yield more efficiently and accurate estimates and enables construction of GW template banks for CBCs, which are then used to detect and extract signals from detector data (e.g., Owen 1996; Magee et al. 2019; Cannon et al. 2020; Luo et al. 2020; Amaro-Seoane et al. 2023; Ren et al. 2023; The LIGO Scientific Collaboration et al. 2025d; Joshi et al. 2025a, b; Niu et al. 2025; Luo et al. 2025a). Below, we briefly introduce the generation of GW waveforms.
@@
where $F_{+}$ and $F_{\times}$ are the antenna pattern function of the GW detector. In practical GW data analysis, we usually Fourier-transform the signal to the frequency domain for convenience, which is equivalent to working in the time-domain. An example for the GW waveform can be given by the first direct detection of GWs, GW150914, with component masses of $\sim 36M_{\odot}$ and $\sim 29M_{\odot}$ at a luminosity distance of $\sim 410$ Mpc, its strain amplitude at Earth is of the order $10^{-21}$, showcasing the remarkable measurement capabilities required for the GW detection (Abbott et al. 2016b).
@@
The above expressions are only approximations under the Newtonian limit, which cannot be applied to late time of inspiral, merger, and ring-down stages under extremely strong gravitational field. Generally speaking, the waveform in the later inspiral stage is often calculated via the post-Newtonian expansion (e.g., Thorne 1980; Damour et al. 2000; Pretorius 2023; Blanchet 2014), in which the higher order nonlinear terms of the field equation are considered. Contributions scaling as $(v/c)^{2n}$ are labeled as the $n\rm PN$ term in the post-Newtonian expansion. Extending the series to high-order PN term is time-consuming and tedious, and to date waveform models consider such terms only up to $4.5\rm PN$ in the literature (Messina & Nagar 2017). As for the merger process, the GW waveform may be calculated by using numerical relativity (e.g., Baumgarte & Shapiro 2010) through $3+1$ decomposition adopting the Arnowitt-Deser-Misner (ADM) formalism (Arnowitt et al. 2008), in which the spacetime is split with several space-like hyperspace. When it comes to the ring-down process, the GW waveform can be modeled with the BH perturbation theory (e.g., Ajith et al. 2008), by solving the Teukolsky and Zerilli equations (Teukolsky 1972; Price 1972). Here we do not intend to overview the details of the GW calculations for these stages, which can be seen in the classical textbook of Maggiore (2007) and Maggiore (2018) for interested readers. However, we noticed that it is not possible to calculate the waveform across all the allowed parameter space, as GW waveform calculations by the above methods are indeed time-consuming. Therefore, in practical searches and data analysis, the GW template is normally estimated by data-driven phenomenological waveform model, which balance computational efficiency with physical fidelity. These models parameterize the signal with a small number of physical parameters (e.g., masses, spins), enabling simultaneous parameter estimation. There are three main phenomenological model series, i.e., IMRPhenom Series (e.g., Pürrer et al. 2023; Khan et al. 2019; Dietrich et al. 2017), SEOBNR Series (e.g., Ramos-Buades et al. 2023; Kumar et al. 2015), and NRSur Models (e.g., Islam et al. 2023; Gramaxo Freitas et al. 2024), which have been widely used in GW data analysis. Currently, these models have been integrated within the main GW data analysis tool package PyCBC  [^2] https://pycbc.org/ (Biwer et al. 2019) and readers can generate GW waveforms for various sources with various parameters conveniently via this package.
@@ 2.1.2 GW Data Analysis
where $A_{m}=e^{-2\pi if((\hat{r}_{m}-\hat{r}_{1})\cdot\hat{n}_{\rm GW})}$ is the phase transfer function, $\hat{r}_{m}$ is the location of the $m$-th detector, $\hat{n}_{\rm GW}$ is the unit direction vector of the GW source, and $S_{{\rm n},m}$ denotes the one-sided power spectrum of the corresponding $m$-th GW detector (e.g., Wen & Chen 2010). Then the optimal squared S/N is given by
@@
where $\Delta\vartheta_{\rm s}$ and $\Delta\phi_{\rm s}$ are the standard deviations obtained from the covariance matrix. As a straight forward magnitude estimation, the median localization precision of sBBH mergers detected by the third-generation GW detectors (such as ET and CE) will be about $\sim 1{\rm deg^{2}}$, which is still large for EM telescope follow-up. More detailed descriptions on the estimation of the GW localization precision for CBCs can be seen in Zhao & Wen (2018) and Pieroni et al. (2022a). In the Fisher information matrix approach we neglect nonlinear signal terms (Vallisneri 2008), an approximation that holds only when the S/N is sufficiently high. It has been shown that for systems with sufficiently high S/Ns ($\geq 10$), parameter estimates from the Fisher information matrix method agree well with those from full Bayesian analysis (e.g., Rodriguez et al. 2013; Grover et al. 2014). However, one should be cautious when applying the Fisher information matrix method to GW events with S/N substantially less than $10$.
@@ 2.1.3 Current and Future GW detectors
    High-frequency ($\rm 10-100$ Hz): The main sources in this band are CBCs (i.e., sBBH, BHNS, and BNS mergers), core-collapse supernovae, and rotating asymmetric NSs. The corresponding detectors are ground-based laser interferometer detectors, including LIGO (e.g., Abbott et al. 2016a), Virgo (Bradaschia et al. 1990), KAGRA (Kagra Collaboration et al. 2019), Indigo (Unnikrishnan 2023), and the third-generation GW detectors, such as CE (Reitze et al. 2019) and ET (Hild et al. 2011).
@@
    Very low-frequency ($\rm 10^{-9}-10^{-7}$ Hz): The loud astrophysical GW sources in this band are inspiraling SMBBHs with masses $\gtrsim 10^{8}M_{\odot}$. Some physical processes occurred in the early universe can also produce stochastic GW background in this band (see Ellis et al. 2024, for a thorough comparison), such as cosmic (super) strings (Hindmarsh & Kibble 1995), cosmological first-order phase transition (Witten 1984) and primodial scalar-induced GW (Tomita 1975), etc. The corresponding detectors are pulsar time arrays (PTAs), including Chinese PTA (CPTA), Parkes PTA (PPTA), North American Nanohertz Observatory for Gravitational waves (NANOGrav), European PTA (EPTA, with Indian PTA, denoted as InPTA), Meerkat PTA (MPTA), and future Square Kilometer Array PTA (SKA-PTA) and next generation Very Large Array (ngVLA) (e.g., Tarafdar et al. 2022; Agazie et al. 2023a; EPTA Collaboration et al. 2023; Reardon et al. 2023; Xu et al. 2023; Miles et al. 2025).
@@ 2.2.1 Geometric Optics
which is similar to that in the standard three-dimensional gravity potential theory. Here we refer the readers to several popular softwares useful for solving the above lens equations numerically for various lens models such as Lenstronomy  [^3] https://lenstronomy.readthedocs.io/en/latest/ (Birrer & Amara 2018), PyAutoLens  [^4] https://pyautolens.readthedocs.io/en/latest/ (Nightingale & Dye 2015) and Lenstool  [^5] https://projets.lam.fr/projects/lenstool/wiki (Kneib et al. 2011).

### 16 May 2026 v2

@@ 1 Introduction
Detection of the gravitational wave (GW) signals from inspiralling and mergers of compact objects [including black holes (BHs) and neutron stars (NSs)] provide a new way to measure their properties and study their origin and nature. The first direct detection of a stellar binary black hole (sBBH) merger (GW150914) by the Laser Interferometer Gravitational wave Observatory (LIGO) confirmed the existence of sBBHs and substantially revised our understanding of sBBH formation (e.g., Abbott et al. 2016b, 2019a, 2020b; The LIGO Scientific Collaboration et al. 2021a, b). The subsequent observations of the binary neutron star (BNS) merger event, GW170817, including its multiwavelength electromagnetic (EM) counterparts not only confirmed the predictions about short $\gamma$-ray burst (sGRB) with afterglow signals and kilonova produced by BNS mergers, but also marked the opening of the new era of multi-messenger astronomy (e.g., Abbott et al. 2017b, a; Coulter et al. 2017; Pian et al. 2017; Dobie et al. 2018; Coughlin et al. 2019; Abbott et al. 2019b). With the multimessenger (GW + EM) information, one may estimate the mass of the tidal debris of BNS by the observation of the kilonova signals and therefore further constrain the fundamental physics related to the equation of state (EOS) of neutron stars (NSs) (e.g., Most et al. 2018; De et al. 2018; Gill et al. 2019). Moreover, with the redshift measurement from the EM signals and the luminosity distance inferred from GW, one can also constrain the Hubble constant to a precision of $\sim 10\%$ with this single object (GW170817) by the “standard siren” method (e.g., Schutz 1986; Abbott et al. 2017c; Soares-Santos et al. 2019). These results underscore the strong prospects for observing compact-binary coalescences (CBCs) via GWs and EMs.
@@
The astrophysical GW sources are quite abundant in the universe and an extremely large number of such sources are expected to be detected in future. Since the first direct detection of GWs by LIGO, more than $300$ stellar CBCs have been detected by LIGO and Virgo, including about three hundreds sBBH mergers, two BNS mergers, and several BHNS mergers (e.g., The LIGO Scientific Collaboration et al. 2025c, a)111see also at the website https://gracedb.ligo.org/. With further upgrade, the Advanced LIGO plus (LIGO A+) (Abbott et al. 2020a) and LIGO Voyager (Adhikari et al. 2020) will detect many more in the near future. The third-generation ground-based GW detectors, i.e., Einstein Telescope (ET) (Hild et al. 2011) and Cosmic Explorer (CE) (Reitze et al. 2019) are even anticipated to detect more than several tens of thousands of CBCs per year. In addition, the pulsar timing arrays (PTA) (e.g., Agazie et al. 2023a; EPTA Collaboration et al. 2023; Reardon et al. 2023; Xu et al. 2023; Miles et al. 2025) are expected to detect the nanohertz GW signals from inspiralling supermassive binary black holes (SMBBHs) with masses $\gtrsim 10^{8}M_{\odot}$ (e.g., Sesana 2013a; Chen et al. 2024a) both collectively as a stochastic background and individually as continuous gravitational wave sources (CGWs). The space-based GW detectors, such as Laser Interferometer Space Antenna (LISA) (Amaro-Seoane et al. 2017), Taiji (Luo et al. 2020) and Tianqin (Mei et al. 2021), are designed to detect the millihertz GW signals from mergers of massive binary black holes (MBBHs) with masses $\sim 10^{4}-10^{7}M_{\odot}$ and extreme/intermediate mass-ratio inspirals (EMRIs/IMRIs) (e.g., Amaro-Seoane et al. 2007; Babak et al. 2017). The numbers of such sources in the millihertz and nanohertz bands that may be detected in future can also be substantial.
@@
A small fraction of various types of GW events may be gravitationally lensed by intervening objects/systems, such as stars, galaxies, and galaxy clusters, and they may be detectable and identifiable by future GW detectors. Since the GWs considered for detection (by ground-based GW detectors, space-based GW detectors, and the PTAs) have much longer wavelengths ($\lambda_{\rm GW}$) comparing with those of EM waves from radio to gamma-ray bands, the wave nature of GWs may have to be considered in some cases, especially when $\lambda_{\rm GW}\gg\lambda_{\rm E}$ with $\lambda_{\rm E}$ representing the Einstein radius of the lens (a characteristic scale). In those cases with $\lambda_{\rm GW}\ll\lambda_{\rm E}$, the gravitational lensing of GWs is in the geometric-optics regime, and the lensing effect is similar to that for EM waves (such as lensed QSOs, lensed galaxies, lensed planets, etc.). However, in some cases the GW wavelength can be larger than the Einstein radius of the lens, i.e., $\lambda_{\rm GW}\gtrsim\lambda_{\rm E}$, thus the wave-optics should apply for the lensing and the diffraction is important, which hardly occur for the cases of EM wave lensing. In addition, not like the cases for EM waves that may be easily contaminated by light from foreground galaxies or distorted by interstellar medium (ISM), the GW signals are hardly contaminated and thus encode clean information about physical properties of the sources and the propagation path. Therefore, the lensing of GW signals offers unique a probe for some key astrophysical and cosmological questions (e.g., Grespan and Biesiada 2023; Smith et al. 2025; More and Phurailatpam 2025; Levan et al. 2025), including the wave nature of GWs, the nature of the dark matter (e.g., Guo and Lu 2022; Fairbairn et al. 2023), the origin of sBBH mergers (e.g., Chen et al. 2022b), and the cosmological parameters at relatively high redshifts (Liao et al. 2017b; Hannuksela et al. 2020; Chen et al. 2025b; Shan et al. 2023).
@@ 2.1.1 GW Waveform
The GW waveform is different in different evolution stages of binary mergers, i.e., inspiral, merger, and ring-down. In the merger and ring-down stages, the waveform can only be accurately calculated by using numerical relativity, for their extremely complex spacetime structure, which is computationally expensive. In the inspiral stage, one can analytically obtain the waveform, and such a method also enables an order of magnitude estimation for the waveform in the other two stages (Bondi 1960; Bondi et al. 1962). Combining the analytical and numerical methods yield more efficiently and accurate estimates and enables construction of GW template banks for CBCs, which are then used to detect and extract signals from detector data (e.g., Owen 1996; Magee and others 2019; Cannon and others 2020; Luo et al. 2020; Amaro-Seoane et al. 2023; Ren et al. 2023; The LIGO Scientific Collaboration et al. 2025d; Joshi and others 2025a, b; Niu et al. 2025; Luo et al. 2025). Below, we briefly introduce the generation of GW waveforms.
@@
where $F_{+}$ and $F_{\times}$ are the antenna pattern function of the GW detector. In practical GW data analysis, we usually Fourier-transform the signal to the frequency domain for convenience, which is equivalent to working in the time-domain. An example for the GW waveform can be given by the first direct detection of GWs, GW150914, with component masses of $\sim 36M_{\odot}$ and $\sim 29M_{\odot}$ at a luminosity distance of $\sim 410$ Mpc, its strain amplitude at Earth is of the order $10^{-21}$, showcasing the remarkable measurement capabilities required for the GW detection (Abbott et al. 2016a).
@@
The above expressions are only approximations under the Newtonian limit, which cannot be applied to late time of inspiral, merger, and ring-down stages under extremely strong gravitational field. Generally speaking, the waveform in the later inspiral stage is often calculated via the post-Newtonian expansion (e.g., Thorne 1980; Damour et al. 2000; Pretorius 2023; Blanchet 2014), in which the higher order nonlinear terms of the field equation are considered. Contributions scaling as $(v/c)^{2n}$ are labeled as the $n\rm PN$ term in the post-Newtonian expansion. Extending the series to high-order PN term is time-consuming and tedious, and to date waveform models consider such terms only up to $4.5\rm PN$ in the literature (Messina and Nagar 2017). As for the merger process, the GW waveform may be calculated by using numerical relativity (e.g., Baumgarte and Shapiro 2010) through $3+1$ decomposition adopting the Arnowitt-Deser-Misner (ADM) formalism (Arnowitt et al. 2008), in which the spacetime is split with several space-like hyperspace. When it comes to the ring-down process, the GW waveform can be modeled with the BH perturbation theory (e.g., Ajith et al. 2008), by solving the Teukolsky and Zerilli equations (Teukolsky 1972; Price 1972). Here we do not intend to overview the details of the GW calculations for these stages, which can be seen in the classical textbook of Maggiore (2007) and Maggiore (2018) for interested readers. However, we noticed that it is not possible to calculate the waveform across all the allowed parameter space, as GW waveform calculations by the above methods are indeed time-consuming. Therefore, in practical searches and data analysis, the GW template is normally estimated by data-driven phenomenological waveform model, which balance computational efficiency with physical fidelity. These models parameterize the signal with a small number of physical parameters (e.g., masses, spins), enabling simultaneous parameter estimation. There are three main phenomenological model series, i.e., IMRPhenom Series (e.g., Pürrer et al. 2023; Khan et al. 2019; Dietrich et al. 2017), SEOBNR Series (e.g., Ramos-Buades et al. 2023; Kumar et al. 2015), and NRSur Models (e.g., Islam et al. 2023; Gramaxo Freitas et al. 2024), which have been widely used in GW data analysis. Currently, these models have been integrated within the main GW data analysis tool package PyCBC  [^2] https://pycbc.org/ (Biwer et al. 2019) and readers can generate GW waveforms for various sources with various parameters conveniently via this package.
@@ 2.1.2 GW Data Analysis
where $A_{m}=e^{-2\pi if((\hat{r}_{m}-\hat{r}_{1})\cdot\hat{n}_{\rm GW})}$ is the phase transfer function, $\hat{r}_{m}$ is the location of the $m$-th detector, $\hat{n}_{\rm GW}$ is the unit direction vector of the GW source, and $S_{{\rm n},m}$ denotes the one-sided power spectrum of the corresponding $m$-th GW detector (e.g., Wen and Chen 2010). Then the optimal squared S/N is given by
@@
where $\Delta\vartheta_{\rm s}$ and $\Delta\phi_{\rm s}$ are the standard deviations obtained from the covariance matrix. As a straight forward magnitude estimation, the median localization precision of sBBH mergers detected by the third-generation GW detectors (such as ET and CE) will be about $\sim 1{\rm deg^{2}}$, which is still large for EM telescope follow-up. More detailed descriptions on the estimation of the GW localization precision for CBCs can be seen in Zhao and Wen (2018) and Pieroni et al. (2022a). In the Fisher information matrix approach we neglect nonlinear signal terms (Vallisneri 2008), an approximation that holds only when the S/N is sufficiently high. It has been shown that for systems with sufficiently high S/Ns ($\geq 10$), parameter estimates from the Fisher information matrix method agree well with those from full Bayesian analysis (e.g., Rodriguez et al. 2013; Grover et al. 2014). However, one should be cautious when applying the Fisher information matrix method to GW events with S/N substantially less than $10$.
@@ 2.1.3 Current and Future GW detectors
    High-frequency ($\rm 10-100$ Hz): The main sources in this band are CBCs (i.e., sBBH, BHNS, and BNS mergers), core-collapse supernovae, and rotating asymmetric NSs. The corresponding detectors are ground-based laser interferometer detectors, including LIGO (e.g., Abbott et al. 2016b), Virgo (Bradaschia et al. 1990), KAGRA (Kagra Collaboration et al. 2019), Indigo (Unnikrishnan 2023), and the third-generation GW detectors, such as CE (Reitze et al. 2019) and ET (Hild et al. 2011).
@@
    Very low-frequency ($\rm 10^{-9}-10^{-7}$ Hz): The loud astrophysical GW sources in this band are inspiraling SMBBHs with masses $\gtrsim 10^{8}M_{\odot}$. Some physical processes occurred in the early universe can also produce stochastic GW background in this band (see Ellis et al. 2024, for a thorough comparison), such as cosmic (super) strings (Hindmarsh and Kibble 1995), cosmological first-order phase transition (Witten 1984) and primodial scalar-induced GW (Tomita 1975), etc. The corresponding detectors are pulsar time arrays (PTAs), including Chinese PTA (CPTA), Parkes PTA (PPTA), North American Nanohertz Observatory for Gravitational waves (NANOGrav), European PTA (EPTA, with Indian PTA, denoted as InPTA), Meerkat PTA (MPTA), and future Square Kilometer Array PTA (SKA-PTA) and next generation Very Large Array (ngVLA) (e.g., Tarafdar et al. 2022; Agazie et al. 2023a; EPTA Collaboration et al. 2023; Reardon et al. 2023; Xu et al. 2023; Miles et al. 2025).
@@ 2.2.1 Geometric Optics
which is similar to that in the standard three-dimensional gravity potential theory. Here we refer the readers to several popular softwares useful for solving the above lens equations numerically for various lens models such as Lenstronomy  [^3] https://lenstronomy.readthedocs.io/en/latest/ (Birrer and Amara 2018), PyAutoLens  [^4] https://pyautolens.readthedocs.io/en/latest/ (Nightingale and Dye 2015) and Lenstool  [^5] https://projets.lam.fr/projects/lenstool/wiki (Kneib et al. 2011).